OpenStudy (anonymous):

4(-2d-3)=12 help plz!

OpenStudy (anonymous):

You want to find d? What is the question?

OpenStudy (anonymous):

4(-2d-3)=12 Thats the question lol

OpenStudy (anonymous):

So I guess you want to find d then... 4(-2d-3)=12 -2d-3=3 -2d=6 -d=3 d=-3